OpenStudy (anonymous):

need help solving this. The value of a building bought in 1990 increases as time passes. Seven years after the building was bought, it was worth $165,000;12 years after it was bought it was worth $180,000. if this relationship between number of years past 1990 and value of building is linear, write an equation describing this relationship. Use ordered pairs of the form (years past 1990, value of building.

OpenStudy (anonymous):

y = mx + b. So what's your slope? Let the x-axis be years and the y-axis be worth.

OpenStudy (anonymous):

Or you can use point-slope form: y - y' = m (x - x'), where (x' , y') is a point on the line.

OpenStudy (anonymous):

first you need your slope, which is -3000/year, then use point slope form: y-180,000=-3,000(x-12) y-180,000=-3,000x+36,000 y=-3,000x+216,000